OpenStudy (anonymous):

Matt is making cookies. He adds 3 over 4 cup of sugar to a bowl. He then removes two 1 over 8 cups of sugar. Which of the following best explains the amount of sugar left in the bowl?

OpenStudy (anonymous):

tal tal the multiple choices are too long and i cant copy and paste sorry

OpenStudy (tal_tal):

Well you subtract 3/4 and 1/8 do you know how?

OpenStudy (anonymous):

nope

OpenStudy (tal_tal):

okay let me try to explain this hold on

OpenStudy (anonymous):

no prob

OpenStudy (tal_tal):

1/8 = 1/8 - you're set for the first one. 3/4 x 2/2 = 6/8 - all I did there was multiply the fraction by '1' only in the form of '2/2.' 1/8 - 6/8 Just subtract the numerators. -5/8 -- ANSWER. Hope this helped, and good luck!! :)

OpenStudy (anonymous):

that is not one of my multiple choices so idk.....

OpenStudy (tal_tal):

Oh can you just put the numbers for your choices

OpenStudy (anonymous):

3/8,3/8, 4/8, 4/8

OpenStudy (mathstudent55):

@Tal_Tal An important part of solving a math problem is getting an answer that makes sense. The problem states that a person added an amount of sugar to a bowl. Then some sugar was removed from the bowl. You can end up with some sugar or no sugar in the bowl, but your answer of a negative amount makes no sense because it has no meaning in the real world. That means your answer cannot be correct.

OpenStudy (mathstudent55):

@abby1012 I'm trying to understand the problem. At first, 3/4 cup of sugar was added to the bowl, right?

OpenStudy (tal_tal):

Thank you for pointing that out I'll keep that in mind next time!

OpenStudy (anonymous):

um okay so if his own is wrong what is it ?

OpenStudy (anonymous):

yes @mathstudent55 but it has because and then it has numbers sorry cant copy and paste

OpenStudy (mathstudent55):

3/4 cup was added to the bowl. The 2 amounts of 1/8 cup each were taken from the bowl. Is this what the problem is telling you?

OpenStudy (mathstudent55):

@abby1012 It's ok. That's why I'm asking to make sure I understand the problem before I try to help you.

OpenStudy (anonymous):

yes i guess

OpenStudy (anonymous):

Matt is making cookies. He adds 3 over 4 cup of sugar to a bowl. He then removes two 1 over 8 cups of sugar. Which of the following best explains the amount of sugar left in the bowl?

OpenStudy (anonymous):

that is all i can copy and paste

OpenStudy (mathstudent55):

\(\sf He ~adds ~3/4 ~cup ~of ~sugar ~to ~a ~bowl.\) \(\sf He ~then ~removes ~\color{red}{two}~ 1/8 ~cups ~of ~sugar.\)

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

yip

OpenStudy (mathstudent55):